# Empirical Analysis of Transaction Conflicts in Ethereum and Solana for Parallel Execution

### v1

@@ Abstract
This paper presents a comprehensive analysis of historical data across two popular blockchain networks: Ethereum and Solana. Our study focuses on two key aspects: transaction conflicts and the maximum theoretical parallelism within historical blocks. We aim to quantify the degree of transaction parallelism and assess how effectively it can be exploited by systematically examining block-level characteristics, both within individual blocks and across different historical periods. In particular, this study is the first of its kind to leverage historical transactional workloads to evaluate transactional conflict patterns. By offering a structured approach to analyzing these conflicts, our research provides valuable insights and an empirical basis for developing more efficient parallel execution techniques in the Ethereum and Solana Virtual Machines. Our empirical analysis reveals that Ethereum blocks frequently achieve high independence--over 50% in more than 50% of blocks, while Solana blocks contain longer conflict chains, comprising $\sim$59of the block size compared to $\sim$18in Ethereum, reflecting fundamentally different parallel execution dynamics.
@@ 1 Introduction
When block transactions are executed, they may modify the same account address or storage slot in a smart contract, potentially resulting in conflicts if one transaction accesses data modified by another. A conflict occurs when two or more transactions access the same state and at least one of them performs an update operation. The transactions in the block must be executed sequentially in preset order, the order of the transactions in the block, or they must be serializable to the preset order if executed in parallel, to provide deterministic execution across the distributed network. The order in which conflicting transactions are executed directly affects the final state. Transactions are independent if they do not modify the same state, and the outcome of one does not impact the other. The independent transactions can be executed in any order or in parallel, as their execution does not interfere with other transactions.
@@
As such, transaction conflicts can be detected through smart contract bytecode analysis, read-write set analysis, and optimistic execution at runtime. Efficient conflict identification and resolution with minimal overhead, including abort and re-execution costs, are crucial for enhancing performance. The timing and cost of conflict detection and resolution significantly affect the efficiency of parallel execution. The distribution of conflicts and state access patterns within historical blocks is the main focus of this study, along with the metrics we discussed above that directly or indirectly affect the parallel execution of transactions and overall throughput. These insights support the development of techniques that improve execution efficiency and enhance parallelization across blockchain networks.
@@ 2 Motivation
For example, Figure 1 illustrates Ethereum’s read-write oblivious execution and Solana’s read-write aware execution using a block $B_{i}$ containing three transfer transactions. In the existing Ethereum model (Figure 1.(a)), transactions $T_{1},\ T_{2},\ \text{and},\ T_{3}$ are executed sequentially in the preset order. Any parallel execution of these transactions in the EVM requires optimistic execution because the read and write sets are unknown prior to execution, making conflict detection possible only at runtime. Consequently, if conflicting transactions are executed speculatively in parallel, they must be validated and potentially rolled back to satisfy the preset serialization, adding significant execution overhead. For example, if $T_{3}$ executes before $T_{2}$, it must be aborted and re-executed after $T_{2}$ commits (tryC2). In contrast, in the Solana model (Figure 1.(b)) transaction’s read and write sets are known in advance before execution. This allows the runtime to detect that $T_{2}$ and $T_{3}$ conflict on $X_{3}$ and $X_{4}$, while $T_{1}$ is independent. As a result, read-write aware execution can safely execute $T_{1}$ in parallel with $T_{2}$ and defer $T_{3}$, demonstrating how this model allows for efficient and conflict-free parallel execution.
@@
  • Quantify the proportion of independent transactions versus conflict families.
  • Understand the depth and width of conflict chains.
  • Compare the parallelization ceiling under different execution models: read-write aware of Solana vs. read-write oblivious execution of Ethereum.
  • Identify design trade-offs for block transactional execution in current VMs, specifically EVM vs. SVM.
@@ 3 System Model
$\mathcal{SVM}=(\mathcal{S}_{S},\mathcal{B}_{S},\mathcal{E}_{S},\mathcal{R}_{S},\mathcal{D})$ (5)
@@
  • $\mathcal{B}_{S}$ is the set of blocks, each containing a list of transactions $\mathcal{T}_{S}$.
  • $\mathcal{E}_{S}:\mathcal{S}_{S}\times\mathcal{B}_{S}\to\mathcal{S}_{S}$ is the execution function.
@@
  • $\mathcal{D}:\mathcal{T}_{S}\to 2^{\mathcal{S}_{S}}\times 2^{\mathcal{S}_{S}}$ is the function that extracts the read and write sets of a transaction.
@@
This ensures that transactions in a block are only executed concurrently if they do not interfere with each other’s read and write dependencies.
@@ 3. Longest chain of conflicting transactions:
The longest conflict chain represents the longest sequence of transactions where each transaction conflicts with the next in the preset order. It quantifies minimal execution time under parallelism and bounds the theoretical maximum speed-up based on the length of sequential dependencies. Even with the number of available resources (threads/cores), the algorithm must execute the transactions in the chain sequentially because of their interdependencies. The longer the chain, the lower the maximum achievable speedup.
To compute the longest chain, construct a conflict graph $G=(V,E)$, where each vertex $v_{i}\in V$ represents a transaction $T_{i}$, and there is an edge $(v_{i},v_{j})\in E$ if $T_{i}$ and $T_{j}$ conflict. Then, the longest chain of conflicting transactions is:
@@ 6. Total and Write-write conflicts:
Ethereum: In our analysis, we segregate Ethereum transactions into two types: ETH transfer transactions and smart contract transactions, to analyze and understand the access and conflict patterns. ETH transfer transactions perform pure value transfers between externally owned addresses (EOAs) or to smart contract addresses. On the other hand, smart contract transactions interact with sender addresses and contract address(es) to modify blockchain state via function calls and storage updates within the contract(s).
@@
Solana: In contrast to Ethereum, Solana transactions are made up of the account-access specification, a list of accounts to read (read-set) or write (write-set[16]. This specification is added to the transactions upfront by the clients through RPC node interaction. The success or failure of a transaction depends on the freshness of its read-write sets, from the moment it is added to the transaction by the client until its execution at the validator node. Two Solana transactions $T_{i}$ and $T_{j}$ are in conflict if both of them access a common account and at least one of them performs a write operation, as defined earlier in Equation 6. The read-set and write-set simplify our analysis and improve the accuracy of conflict detection for Solana blocks. We used the beta API of Solana’s mainnet to obtain block details in JSON format with the max-supported transaction version set to 0 [17]. The extracted details are then parsed to obtain the information required for our analysis.
@@ 5 Ethereum: An Empirical Study
Block SizeETH TransferSmart Contract TxsERC20 TransferIndependent TxsLongest Chain of ConflictsIndependent ETH TxsIndependent SC TxsConflict Families0050505050100100100100150150150150200200200200
@@ 6 Solana: An Empirical Study
As illustrated in Figure 4 and Table 3, the following observations can be made for real-world Ethereum data:
@@ 7 Conflicts on Ethereum versus Solana
Block SizeIndependent TxnsLongest Chain of ConflictsConflict FamiliesDense Conflict Family Size00100100100100200200200200300300300300400400400400170170170170868686863131313172727272919191913523523523522929292910910910910948484848176176176176Transaction CountEthereumSolana
@@ 8 Concluding Remarks
Appendix A Appendix
Table 4 shows the analysis data of recent Ethereum blocks: 21631001–21631020, we observe a moderate degree of transactional independence and conflict granularity. The average percentage of independent contract transactions is $\sim$45.26%, with some blocks exceeding 50%, indicating substantial room for parallel execution. The longest conflict chains remain relatively shallow, averaging $\sim$32 transactions or 18% of block size, compared to the longer conflict chains seen in Solana. However, certain blocks still show dense conflict family concentration, with the most dominant family (Block 21631001) comprising up to 53.41% of the block size (337). The most accessed EOAs and contracts (e.g., 0x4350...2dd and 0xc02a...c2) appear recurrently, but their access counts are more evenly distributed compared to Solana. In general, the Ethereum execution pattern shows a mix of high independence and scattered conflict clusters, which could result in moderate parallelism under read-write oblivious execution.
Block Number 21631001 21631002 21631003 21631004 21631005 21631006 21631007 21631008 21631009 21631010
Block Size 337 148 82 191 233 154 192 163 177 207
ETH Transfer Txs 130 39 27 65 75 64 67 60 68 86
Smart Contract (SC) Txs 207 109 55 126 158 90 125 103 109 121
SC Deployment Txs 0 0 1 1 1 1 1 0 0 1
ERC20 Transfer Txs 75 46 35 59 55 34 46 47 55 52
Total Conflicts 191 79 35 99 108 68 91 85 96 126
Independent Txs 146 69 47 92 125 86 101 78 81 81
Longest Conflict Chain 56 28 24 50 52 20 28 27 39 52
Independent ETH Txs 82 34 23 53 69 47 57 46 53 37
Independent SC Txs 66 38 27 42 57 40 46 34 31 45
Conflict Families 116 51 38 73 99 74 75 67 70 68
Dense Conf Family Size 180 89 39 106 127 61 111 86 101 101
Most Accessed EOA 0x435…d 0x435…d 0x463…8 0x435…d 0x435…d 0xb5d…1 0x435…d 0x435…d 0x2b3…b 0x924…f
Most Accessed EOA Count 22 15 8 23 25 11 12 20 26 34
Cumulative EOA Accesses 710 318 145 396 554 332 396 333 338 480
Most Accessed SC 0xc02…2 0xc02…2 0xdac…7 0xc02…2 0xc02…2 0xc02…2 0xc02…2 0xdac…7 0xdac…7 0xdac…7
Most Accessed SC Count 72 29 22 37 50 26 47 28 40 35
Cumulative SC Accesses 576 253 94 343 420 236 307 255 268 294
ETH Transfer Txs (%) 38.58 26.35 32.93 34.03 32.19 41.56 34.90 36.81 38.42 41.54
Independent Txs (%) 43.32 46.62 57.32 48.17 53.65 55.84 52.60 47.85 45.76 39.13
Longest Conflict Chain (%) 16.62 18.92 29.27 26.18 22.32 12.99 14.58 16.56 22.03 25.12
Independent ETH Txs (%) 63.08 87.18 85.19 81.54 92.00 73.44 85.07 76.67 77.94 43.02
Independent SC Txs (%) 31.88 34.86 49.09 33.33 36.08 44.44 36.80 33.01 28.44 37.19
Block Number 21631011 21631012 21631013 21631014 21631015 21631016 21631017 21631018 21631019 21631020
Block Size 148 134 175 200 138 180 119 230 145 166
ETH Transfer Txs 46 46 51 66 42 68 38 100 47 55
Smart Contract (SC) Txs 102 88 124 134 96 112 81 130 98 111
SC Deployment Txs 0 1 0 1 0 0 0 2 1 1
ERC20 Transfer Txs 31 32 58 41 40 58 43 46 30 39
Total Conflicts 70 51 82 91 56 104 49 117 61 82
Independent Txs 78 83 93 109 82 76 70 113 84 84
Longest Conflict Chain 20 16 34 21 22 34 26 41 19 26
Independent ETH Txs 20 16 34 55 41 43 34 62 36 45
Independent SC Txs 39 40 45 55 42 33 38 52 48 41
Conflict Families 61 66 72 91 68 61 61 94 67 68
Dense Conf Family Size 87 66 96 85 66 110 50 103 68 87
Most Accessed EOA 0x435…d 0x435…d 0x435…d 0x435…d 0x435…d 0x435…d 0x435…d 0xeba…1 0x435…d 0x435…d
Most Accessed EOA Count 18 14 13 18 15 26 10 18 13 13
Cumulative EOA Accesses 310 303 319 473 314 341 221 509 303 365
Most Accessed SC 0xc02…2 0xc02…2 0xc02…2 0xc02…2 0xc02…2 0xdac…7 0xdac…7 0xdac…7 0xc02…2 0xc02…2
Most Accessed SC Count 41 29 39 38 29 35 27 35 26 41
Cumulative SC Accesses 265 262 299 384 242 226 137 334 251 310
ETH Transfer Txs (%) 31.08 34.33 29.14 33.00 30.43 37.78 31.93 43.48 32.41 33.13
Independent Txs (%) 52.70 61.94 53.14 54.50 59.42 42.22 58.82 49.13 57.93 50.60
Longest Conflict Chain (%) 13.51 11.94 19.43 10.50 15.94 18.89 21.85 17.83 13.10 15.66
Independent ETH Txs (%) 89.13 93.48 96.08 81.82 97.62 63.24 89.47 62.00 76.60 81.82
Independent SC Txs (%) 38.24 45.45 36.29 41.04 43.75 29.46 46.91 40.00 48.98 36.94
Table 5 presents a detailed breakdown of transaction-level conflict metrics across recent Solana blocks (314184230–314184249). The table also reports average and maximum read-write set sizes, which influence conflict likelihood and execution overlap, and includes several additional metrics beyond those discussed in the main draft to provide a more comprehensive view of access patterns and contention behavior. Despite fluctuations in block size and non-voting transaction count, independence remains consistently low, with the percentage of independent non-voting transactions averaging below 5% in most blocks. Write-write conflicts dominate, comprising over 70% of total conflicts on average, and conflict chains often span more than half the block. While the number of conflict families has increased, indicating finer-grained parallelism, the high proportion of conflicting transactions and declining success rates highlight persistent contention challenges under Solana’s read-write aware execution model.
Block Number 314184230 314184231 314184232 314184233 314184234 314184235 314184236 314184237 314184238 314184239
Block Size 2281 1537 2090 1689 1844 1304 2075 1722 1439 2026
NV Txs 590 518 515 335 431 72 446 330 586 517
Voting Txs 1691 1019 1575 1354 1413 1232 1629 1392 853 1509
Successful NV-Txs 454 170 477 298 276 40 403 304 372 377
Failed NV-Txs 136 348 38 37 155 32 43 26 214 140
Largest Rset Size 21 21 26 18 22 20 21 22 21 27
Average Rset Size 6.38 7.88 5.58 5.83 7.98 8.35 6.29 7.19 7.89 7.63
Largest Wset Size 34 27 42 19 34 24 37 35 35 35
Average Wset Size 7.52 10.49 6.43 7.10 9.28 8.76 7.00 8.95 10.27 9.20
Accounts Accessed 2573 1172 2309 1676 2420 437 2326 1861 2816 2623
Most Read Account Comp…1 Comp…1 Comp…1 Comp…1 Comp…1 Comp…1 $1$…1 $1$…1 Comp…1 Comp…1
Most Read Count 449 485 356 255 403 67 323 289 451 403
Most Write Account GZ…ib J3…S7 GZ…ib GZ…ib GZ…ib Ce…iM Ce…iM Ce…iM oQ…d7 Ce…iM
Most Write Count 155 110 95 90 75 16 40 40 44 77
Conflicting NV-Txs 570 496 481 333 428 66 405 319 556 500
Total Conflicts 16612 17116 8955 5903 8619 212 4166 3184 4890 10900
Write-Write Conflicts 16145 17071 8871 5566 5762 212 3803 3184 4739 5906
Longest Conflict Chain 386 326 326 238 226 23 239 232 375 366
Conflict Families 32 33 46 6 6 13 59 20 42 28
Dense Conf Family Size 527 374 407 301 419 29 283 283 519 455
Independent NV-Txs 20 22 34 2 3 6 41 11 30 17
Independent NV-Txs (%) 3.39 4.25 6.60 0.60 0.70 8.33 9.19 3.33 5.12 3.29
NV-Txs (%) 25.87 33.70 24.64 19.83 23.37 5.52 21.49 19.16 40.72 25.52
Successful NV-Txs (%) 76.95 32.82 92.62 88.96 64.04 55.56 90.36 92.12 63.48 72.92
Conflicting NV-Txs (%) 96.61 95.75 93.40 99.40 99.30 91.67 90.81 96.67 94.88 96.71
Conflict Chain (%) 65.42 62.93 63.30 71.04 52.44 31.94 53.59 70.30 63.99 70.79
Block Number 314184240 314184241 314184242 314184243 314184244 314184245 314184246 314184247 314184248 314184249
Block Size 2331 1863 1723 1579 2338 1568 2239 1484 2620 1711
NV Txs 473 498 327 430 644 544 485 477 781 535
Voting Txs 1858 1365 1396 1149 1694 1024 1754 1007 1839 1176
Successful NV-Txs 420 262 188 192 342 348 312 233 308 210
Failed NV-Txs 53 236 139 238 302 196 173 244 473 325
Largest Rset Size 38 24 26 21 24 24 24 24 22 38
Average Rset Size 6.58 8.51 9.38 9.22 8.28 7.93 7.87 8.99 6.70 8.28
Largest Wset Size 38 35 43 50 42 46 46 50 46 46
Average Wset Size 8.10 10.69 13.67 11.23 8.84 10.53 10.28 12.21 6.95 11.59
Accounts Accessed 2649 2283 2814 2191 2906 2933 2629 2779 2730 2492
Most Read Account $1$…1 Comp…1 Comp…1 Comp…1 Comp…1 Comp…1 Comp…1 Comp…1 Comp…1 Comp…1
Most Read Count 368 424 263 380 565 419 394 401 722 510
Most Write Account FR…Qs 7e…Aw HW…oY 4Y…QU Ce…iM Ce…iM VN…nP VN…nP 6r…hE 9D…QX
Most Write Count 44 46 29 61 74 33 40 35 44 84
Conflicting NV-Txs 432 482 311 424 616 525 472 461 755 524
Total Conflicts 6738 5614 2715 7672 12773 5018 11463 8480 15345 9763
Write-Write Conflicts 4073 4715 2102 5887 7727 4332 4036 4009 11245 7219
Longest Conflict Chain 266 304 197 285 299 325 314 309 320 322
Conflict Families 54 22 23 9 38 28 21 19 38 18
Dense Conf Family Size 355 460 295 417 533 490 436 450 394 476
Independent NV-Txs 41 16 16 6 28 19 13 16 26 11
Independent NV-Txs (%) 8.67 3.21 4.89 1.40 4.35 3.49 2.68 3.35 3.33 2.06
NV-Txs (%) 20.29 26.73 18.98 27.23 27.54 34.69 21.66 32.14 29.81 31.27
Successful NV-Txs (%) 88.79 52.61 57.49 44.65 53.11 63.97 64.33 48.85 39.44 39.25
Conflicting NV-Txs (%) 91.33 96.79 95.11 98.60 95.65 96.51 97.32 96.65 96.67 97.94
Conflict Chain (%) 56.24 61.04 60.24 66.28 46.43 59.74 64.74 64.78 40.97 60.19

### v2

@@ Abstract
This paper presents a comprehensive analysis of historical data across two popular blockchain networks: Ethereum and Solana. Our study focuses on two key aspects: transaction conflicts and the maximum theoretical parallelism within historical blocks. We aim to quantify the degree of transaction parallelism and assess how effectively it can be exploited by systematically examining block-level characteristics, both within individual blocks and across different historical periods. In particular, this study is the first of its kind to leverage historical transactional workloads to evaluate transactional conflict patterns. By offering a structured approach to analyzing these conflicts, our research provides valuable insights and an empirical basis for developing more efficient parallel execution techniques for smart contracts in the Ethereum and Solana virtual machines. Our empirical analysis reveals that historical Ethereum blocks frequently achieve high independence, over 50% in more than 50% of blocks, while Solana historical blocks contain longer conflict chains, comprising $\sim$59of the block size compared to $\sim$18in Ethereum, reflecting fundamentally different parallel execution dynamics.
@@ 1 Introduction
When block transactions are executed, they may modify the same account address or storage slot in a smart contract, potentially resulting in conflicts if one transaction accesses data modified by another. A conflict occurs when two or more transactions access the same state and at least one of them performs an update operation. The transactions in the block must be executed sequentially in preset order, the order of the transactions in the block, or they must be serializable to the preset order if executed in parallel, to provide deterministic execution across the distributed network. The order in which conflicting transactions are executed directly affects the final state. Two transactions are independent if they do not modify the same state, and the outcome of one does not impact the other. The independent transactions can be executed in any order or in parallel, as their execution does not interfere with other transactions.
@@
Potential transaction conflicts can be identified through static analysis of smart contract bytecode, while exact conflicts can be detected via optimistic execution by analyzing the read and write sets of each transaction. Efficient conflict identification and resolution with minimal overhead, including abort and re-execution costs, are crucial for enhancing performance. The timing and cost of conflict detection and resolution significantly affect the efficiency of parallel execution. The distribution of conflicts and state access patterns within historical blocks is the main focus of this study, along with the metrics we discussed above that directly or indirectly affect the parallel execution of transactions and overall throughput. These insights support the development of techniques that improve execution efficiency and enhance parallelization across blockchain networks.
@@ 2 Motivation
For example, Figure 1 illustrates Ethereum’s read-write oblivious execution and Solana’s read-write aware execution using a block $B_{i}$ containing three transfer transactions. In the existing Ethereum model (see Figure 1a), transactions $T_{1},\ T_{2},\ \text{and},\ T_{3}$ are executed sequentially in the preset order. Any parallel execution of these transactions in the EVM requires optimistic execution because the read and write sets are unknown prior to execution, making conflict detection possible only at runtime. Consequently, if conflicting transactions are executed speculatively in parallel, they must be validated and potentially rolled back to satisfy the preset serialization, adding significant execution overhead. For example, if $T_{3}$ executes before $T_{2}$, it must be aborted and re-executed after $T_{2}$ commits (tryC2). In contrast, in the Solana model transaction’s read and write sets are known in advance before execution, as shown in Figure 1b. This allows the runtime to detect that $T_{2}$ and $T_{3}$ conflict on $X_{3}$ and $X_{4}$, while $T_{1}$ is independent. As a result, read-write aware execution can safely execute $T_{1}$ in parallel with $T_{2}$ and defer $T_{3}$, demonstrating how this model allows for efficient and conflict-free parallel execution.
@@
RQ1
Quantify what fraction of transactions are independent and how transactions distributed across conflict families?
RQ2
What are the typical depth and width characteristics of conflict chains in historical blocks?
RQ3
How does the maximum achievable parallelism compare between read-write aware execution (as in Solana) and read-write oblivious execution (as in Ethereum)?
RQ4
What are the fundamental design trade-offs in block transactional execution across different virtual machines, particularly the EVM and the SVM?
@@ 3 System Model
$\mathcal{SVM}=(\mathcal{S}_{S},\mathcal{B}_{S},\mathcal{E}_{S},\mathcal{R}_{S})$ (5)
@@
  • $\mathcal{B}_{S}$ is the set of blocks, each containing a list of transactions $\mathcal{T}_{S}$, with associated read sets (Rset) and write sets (Wset).
  • $\mathcal{E}_{S}:\mathcal{S}_{S}\times\mathcal{B}_{S}\to\mathcal{S}_{S}$ is the parallel execution function. Given the global state and a block, it applies a schedule over the transactions $\mathcal{T}_{S}$ using access-set profiling to determine a non-conflicting execution order. Internally, this schedule respects conflicts defined by overlaps in Rset and Wset.
@@
This ensures that transactions in a block are only executed concurrently if they do not interfere with each other’s read and write sets.
@@ 3. Longest chain of conflicting transactions:
The longest conflict chain, or the longest chain of dependent transactions, represents the longest sequence of transactions where each transaction conflicts with the transaction next in the preset order. It quantifies minimal execution time under parallelism and bounds the theoretical maximum speed-up based on the length of sequential dependencies. Even with the number of available resources (threads/cores), the algorithm must execute the transactions in the chain sequentially because of their interdependencies. The longer the chain, the lower the maximum achievable speedup.
To compute the longest chain of conflict-dependent transactions, construct a conflict graph $G=(V,E)$, where each vertex $v_{i}\in V$ represents a transaction $T_{i}$, and there is an edge $(v_{i},v_{j})\in E$ if $T_{i}$ and $T_{j}$ conflict. Then, the longest chain of conflicting transactions is:
@@ 6. Total and Write-write conflicts:
Ethereum
In our analysis, we segregate Ethereum transactions into two types: ETH transfer transactions and smart contract transactions, to analyze and understand the access and conflict patterns. ETH transfer transactions perform pure value transfers between externally owned addresses (EOAs) or to smart contract addresses. On the other hand, smart contract transactions interact with sender addresses and contract address(es) to modify blockchain state via function calls and storage updates within the contract(s).
@@
Solana
In contrast to Ethereum, Solana transactions are made up of the account-access specification, a list of accounts to read (read-set) or write (write-set[16]. This specification is added to the transactions upfront by the clients through RPC node interaction. The success or failure of a transaction depends on the freshness of its read-write sets, from the moment it is added to the transaction by the client until its execution at the validator node. Two Solana transactions $T_{i}$ and $T_{j}$ are in conflict if both of them access a common account and at least one of them performs a write operation, as defined earlier in Equation 6. The read-set and write-set simplify our analysis and improve the accuracy of conflict detection for Solana blocks. We used the beta API of Solana’s mainnet to obtain block details in JSON format with the max-supported transaction version set to 0 [17]. The extracted details are then parsed to obtain the information required for our analysis.
@@ 5 Ethereum: An Empirical Study
Block SizeETH TransferSmart Contract TxsERC20 TransferIndependent TxsLongest Chain of ConflictsIndependent ETH TxsIndependent SC TxsConflict Families050100150200
@@
Table 3: Ethereum recent blocks (21631001–21631020): analysis on additional metrics.
Block ID Block Size SC Deployment Txs Total Conflicts Most Accessed EOA Cumulative EOA Accesses Most Accessed SC Cumulative SC Accesses Chain of Conflicts (%)
Address Access Count Address Access Count
21631001 337 0 191 0x435...d 22 710 0xc02...2 72 576 16.62
21631002 148 0 79 0x435...d 15 318 0xc02...2 29 253 18.92
21631003 82 1 35 0x463...8 8 145 0xdac...7 22 94 29.27
21631004 191 1 99 0x435...d 23 396 0xc02...2 37 343 26.18
21631005 233 1 108 0x435...d 25 554 0xc02...2 50 420 22.32
21631006 154 1 68 0xb5d...1 11 332 0xc02...2 26 236 12.99
21631007 192 1 91 0x435...d 12 396 0xc02...2 47 307 14.58
21631008 163 0 85 0x435...d 20 333 0xdac...7 28 255 16.56
21631009 177 0 96 0x2b3...b 26 338 0xdac...7 40 268 22.03
21631010 207 1 126 0x924...f 34 480 0xdac...7 35 294 25.12
21631011 148 0 70 0x435...d 18 310 0xc02...2 41 265 13.51
21631012 134 1 51 0x435...d 14 303 0xc02...2 29 262 11.94
21631013 175 0 82 0x435...d 13 319 0xc02...2 39 299 19.43
21631014 200 1 91 0x435...d 18 473 0xc02...2 38 384 10.50
21631015 138 0 56 0x435...d 15 314 0xc02...2 29 242 15.94
21631016 180 0 104 0x435...d 26 341 0xdac...7 35 226 18.89
21631017 119 0 49 0x435...d 10 221 0xdac...7 27 137 21.85
21631018 230 2 117 0xeba...1 18 509 0xdac...7 35 334 17.83
21631019 145 1 61 0x435...d 13 303 0xc02...2 26 251 13.10
21631020 166 1 82 0x435...d 13 365 0xc02...2 41 310 15.66
Observation-5:
Table 3 presents additional metrics for block-wise trends of the recent historical blocks 21631001–21631020, as shown in Table 2. These metrics complement the earlier conflict-based analysis in Table 2 by revealing structural dependencies for parallel execution.
From Table 3, we can observe that a small number of EOAs and smart contract accounts are frequently accessed across transactions and become hotspots within each block; e.g., 0x435…d and 0xc02…2 appear as the most accessed addresses in most of the blocks. This indicates a high degree of access skew across different blocks and within each block, which leads to hotspots that can serialize execution in otherwise parallelizable workloads; however, several of these accesses could be just in read mode. For instance, block 21631003 contains only 82 transactions with 145 cumulative EOA accesses, and over 5% of these target a single EOA. This illustrates that even small blocks can exhibit significant access skew. Despite block sizes that vary between 82 and 337 transactions, the average length of the longest conflict chain 18.23% of block size, showing dominant bottlenecks are not due to block size but to contention on shared addresses. Notably, blocks like 21631003 and 21631013 show relatively fewer total transactions; however, they consist of longer dependency chains due to concentrated access to shared smart contracts. These observations suggest that access frequency and locality are stronger indicators of the scope of parallel execution than the raw transaction count.
@@ 6 Solana: An Empirical Study
As illustrated in Figure 4 and Tables 4 and 5, the following observations can be made for real-world Solana blocks:
@@
Table 5: Solana recent historical blocks (314184230–314184249): additional analysis of read-write sets and account access patterns.
Block ID Non-Voting (NV) Txs Largest Rset Size Average Rset Size Largest Wset Size Average Wset Size Accounts Accessed Most Read Account Most Written Account
Address Count Address Count
314184230 590 21 6 34 8 2573 Comp...1 449 GZ...ib 155
314184231 518 21 8 27 10 1172 Comp...1 485 J3...S7 110
314184232 515 26 6 42 6 2309 Comp...1 356 GZ...ib 95
314184233 335 18 6 19 7 1676 Comp...1 255 GZ...ib 90
314184234 431 22 8 34 9 2420 Comp...1 403 GZ...ib 75
314184235 72 20 8 24 9 437 Comp...1 67 Ce...iM 16
314184236 446 21 6 37 7 2326 $1$...1 323 Ce...iM 40
314184237 330 22 7 35 9 1861 $1$...1 289 Ce...iM 40
314184238 586 21 8 35 10 2816 Comp...1 451 oQ...d7 44
314184239 517 27 8 35 9 2623 Comp...1 403 Ce...iM 77
314184240 473 38 7 38 8 2649 $1$...1 368 FR...Qs 44
314184241 498 24 9 35 11 2283 Comp...1 424 7e...Aw 46
314184242 327 26 9 43 14 2814 Comp...1 263 HW...oY 29
314184243 430 21 9 50 11 2191 Comp...1 380 4Y...QU 61
314184244 644 24 8 42 9 2906 Comp...1 565 Ce...iM 74
314184245 544 24 8 46 11 2933 Comp...1 419 Ce...iM 33
314184246 485 24 8 46 10 2629 Comp...1 394 VN...nP 40
314184247 477 24 9 50 12 2779 Comp...1 401 VN...nP 35
314184248 781 22 7 46 7 2730 Comp...1 722 6r...hE 44
314184249 535 38 8 46 12 2492 Comp...1 510 9D...QX 84
Observation-5:
Table 5 presents additional metrics for block-wise trends of recent Solana blocks 314184230–314184249, analyzed through their transaction-level read-write sets. These metrics complement earlier contention or scheduling analyses by characterizing structural features that influence Solana’s parallel execution model.
From Table 5, we observe consistently high reuse of specific accounts across blocks, especially the ComputeBudget1111... account, which is read in 17 out of the 20 blocks and dominates access frequency (e.g., 722 reads in block 314184248). Similarly, certain smart contract accounts like CebN5W... and VN1RtW... recur frequently as the most written accounts. This illustrates a access skewness, where a few accounts become systemic hotspots across blocks. Interestingly, the number of non-voting transactions varies significantly, from 72 (block 314184235) to 781 (block 314184248), however access contention remains prominent regardless of block size. For example, both blocks 314184240 and 314184249 show the largest Rset and Wset sizes (up to 38 and 46 accounts, respectively), reflecting extensive data dependencies. These structural patterns suggest that Solana’s optimistic parallel execution must still grapple with dependency bottlenecks arising from frequent access to shared accounts. Overall, the results indicate that access locality and skew, particularly with respect to protocol-level accounts (e.g., ComputeBudget1111…), for system-level reads or writes play a dominant role in shaping the effective parallelism, often more than block size or average set size alone.
@@ 7 Conflicts on Ethereum versus Solana
Block SizeIndependent TxnsLongest Chain of ConflictsConflict FamiliesDense Conflict Family Size050100150200250300350400170170170170868686863131313172727272919191913523523523522929292910910910910948484848176176176176Transaction CountEthereumSolana